Soal SMA · Soal cerita · Kesulitan 4/5
In a certain city, the probability that it rains on any given day is 0.3. If it rains, the probability that a particular event is cancelled is 0.8. If it does not rain, the probability that the event is cancelled is 0.1. What is the probability that the event is cancelled on a randomly chosen day?
- A.0.24
- B.0.31
- C.0.38
- D.0.45
Kunci jawaban & pembahasan
Jawaban: 0.31
P(cancelled) = P(rain)*P(cancel|rain) + P(no rain)*P(cancel|no rain) = 0.3*0.8 + 0.7*0.1 = 0.24 + 0.07 = 0.31. Thus, option B is correct.
Mau penjelasan yang menyesuaikan level anak?
Xinau AI membahas soal langkah demi langkah, menyesuaikan gaya bahasa dengan usia anak, dan menyusun latihan berikutnya berdasarkan jawaban sebelumnya.
Coba gratis 30 hari